OpenStudy (anonymous):

CAN SOMEBODY PLEASE HELP. -9x + 3y = 6 Its in standard form but we have to put it in slope intercept form and i got y = 3x - 2 is that right..??? And i have to find the vertical intercept (y) and horizontal (x)

OpenStudy (anonymous):

i mean 3x + 2

OpenStudy (anonymous):

yes y=3x+2 is correct for the slope intercept form So in order to find the x and y intercept, you need to set x and y to 0 accordingly so you want to find (x,0) and (0,y) so first to find the y intercept. the y intercept is when x =0 so y=3x+2 y=(3*0)+2 y=2 so x=0, y=2 is your y intercept now to find the x intercept, set y=0 y=3x+2 0=3x+2 -2=3x x=(-2/3) so x intercept is x=-2/3 y=0 Is that clear?

OpenStudy (anonymous):

yes my friends were helping me but i didnt know how they got 3x instead of -3x ..

OpenStudy (anonymous):

oh okay :)

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

No problem!